OpenStudy (anonymous):

Help! Solving Systems of Equations Approximately. Algebra I. Medal, Fan, Testimonial!!!!!

OpenStudy (anonymous):

Part A: Explain why the x-coordinates of the points where the graphs of the equations y = 4^x and y = 2^x + 2 intersect are the solutions of the equation 4^x = 2^x + 2. (4 points) Part B: Make tables to find the solution to 4^x = 2^x + 2. Take the integer values of x between -3 and 3. (4 points) Part C: How can you solve the equation 4^x = 2^x + 2 graphically? (2 points)

OpenStudy (anonymous):

@PatrickStar @StudyGurl14 @mathmale @CGGURUMANJUNATH @ranga @agent0smith

OpenStudy (anonymous):

A) curve 1 is y=4x^2 and curve 2 is y=2x^2+2 When these curves intersect, the point/points of intersection will have coordinates(x,y) This point will lie on both these curves. So basically we solve this set of equations to find the corresponding values of x and y and the answer we get will be their point of intersection

OpenStudy (anonymous):

Oh, okay. So, I think I can do Part C. And is Part B, which is making a table - I just replace "x" with -3, -2, -1, 0, 1, 2, 3, right?

OpenStudy (anonymous):

Yes